OpenStudy (anonymous):

How to solve system of equations 3x+5y=9 and 2x-y=-7

OpenStudy (anonymous):

Kill a black

OpenStudy (anonymous):

you can use the substitution method, addition method, or linear combination method. personally. my favorite is the substitution method.

OpenStudy (anonymous):

yes I would like to use the substitutuin. How can i solve it?

OpenStudy (anonymous):

I got y=-3/5x+9/5 and y=2x+7 where do I substitute?

OpenStudy (anonymous):

nOT GOING TO SOLVE :( DON'T NDERSTAND HEHE:P

OpenStudy (anonymous):

EXPLAIN THE PROBLEM

OpenStudy (anonymous):

Lol are you for real?

OpenStudy (anonymous):

EXPLAIN IT TO ME SO THAT I MAY ASSIST YOU

OpenStudy (anonymous):

I have to solve the sysem of equations 3x+5y=9 and 2x-y= -7

OpenStudy (anonymous):

x=-2 y=3

OpenStudy (anonymous):

what did you substitute? please explain

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Guys, why don't you multiply the 2nd eq with 5 and then add it to 1st equation. In this way, we may get rid of variable "y" and solve for "x". Then put the value of "x" in any of the equations e.g in eq-2, and u get the value of y. And I got x= -2 , y =3

OpenStudy (anonymous):

well lets start with the 2nd equation. 2x-y=7 Since 2x doesn't have a variable to slove for, move it to the other side. -y=2x-7 Multiply everyterm by -1 y=2x+7 Now that you've found your y, place it anywhere you see y in the first equation. 2x+5(2x+7)=9 Multiply 5 by each term in the par. 3x+10x+35=9 Fuse together like terms. 13x+35=9 Subtract 35 from both sides. 13x=-26 Divide and you get x=-2 Now that you have your x value, replace x with -2 in the 2nd equation. y= 2(-2)+7 y=-4+7 y=3 So your solutions are: x= -2 y=3

OpenStudy (anonymous):

3x + 5y = 9 ......eq-1 2x - y = -7 .......eq-2 Multiply e-2 with "5" and add it to the eq-1. 5(2x - y) = 5 (-7) 10x - 5y = -35...now add this to eq-1 3x + 5y = 9 + 10x - 5y = -35 => 3x + 10x + 5y - 5y = 9 - 35 => 13x = -26 => x = -2 Now put the value of "x" into eq-2. 2x - y = -7 2 (-2) - y = -7 -4 - y = -7 -y = -7 + 4 -y = -3 y = 3 So, x = -2 and y = 3 .....Answer

OpenStudy (anonymous):

Thanks a lot mitch and Atif

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

how do I give the medal? lol

OpenStudy (anonymous):

Sorry I am kind of new here.

OpenStudy (anonymous):

you already did :)